OpenStudy (anonymous):

M is the midpoint of . The coordinates of A are (-2,3) and the coordinates of M are (1,0). Find the coordinates of B. --what is the solution on how to get the coordinates . please reply :(

OpenStudy (anonymous):

Use the fact that the midpoint between x1,y1 and x2,y2 is (x1+x2)/2,(y1+y2)/2

OpenStudy (anonymous):

Let (x,y) be the coordinate of B, thus\[\frac{ -2+x }{ 2 }=1\]and\[\frac{ 3+y }{ 2 }=0\] Now will you solve for x and y.

OpenStudy (anonymous):

formula for finding coordinates: B=(x1+x2)/2,(y1+y2 )/2 use this formula

OpenStudy (anonymous):

\[x1+x2/2,y1+y2/2 \] \[x1=-2, y1=3 \]\[x2=1,y2=0\] \[(-2+1/2),(3+0/2)\]

OpenStudy (anonymous):

you can find the further urself =)